A nationwide survey reveals that teenagers spend an average of 3 hours per day on social media with a standard deviation of 1.5 hours. What is the proportion of teenagers who spend more than 5 hours per day on social media? Group of answer choices .0918 .9082 .0951 .0085
Answer #1

Solution:

We are given

µ = 3

σ = 1.5

We have to find P(X>5) = 1 - P(X<5)

Z = (X - µ)/σ

Z = (5 - 3)/1.5

Z = 1.333333

P(Z<1.333333) = P(X<5) = 0.908789

(by using z-table)

P(X>5) = 1 - P(X<5)

P(X>5) = 1 - 0.908789

P(X>5) = 0.0918

Answer: 0.0918